OpenStudy (anonymous):

If y varies directly with x and y = -7 when x = -1, find x when y = -84.

OpenStudy (anonymous):

Do you want explanation too?

OpenStudy (anonymous):

Yeah.

OpenStudy (saifoo.khan):

saifoo.khan to the rescue!

OpenStudy (anonymous):

:D

OpenStudy (saifoo.khan):

ishaan is helping.. =D

OpenStudy (anonymous):

Mmmhmmm :)

OpenStudy (anonymous):

Hmm Okay :/ Y = k *X -7= k* -1 k = 7 Y = k*X -84 = 7*X X = -84/7 = -12 But now that I have given you explanation I would like to see your work too

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Well, I just needed to know how to work the equation.

OpenStudy (anonymous):

Hmm Alright No Problem